How many different plates are there that involve 1, 2 or 3 letters followed by 1,



2, 3 or 4 digits?



How many 2 digit or 3-digit numbers can be formed using the digits 1, 3, 4, 5, 6,



8 and 9 if no repetition is allowed?



1
Expert's answer
2022-01-10T11:37:16-0500

i.


"26(10)+26(10)^2+26(10)^3+26(10)^4"

"+26^2(10)+26^2(10)^2+26^2(10)^3+26^2(10)^4"

"+26^3(10)+26^3(10)^3+26^3(10)^3+26^3(10)^4"

"=218886980"

ii.


"P(7,2)+P(7,3)=\dfrac{7!}{(7-2)!}+\dfrac{7!}{(7-3)!}"

"=42+210=252"